# Coordinate-based Speed of Sound Recovery for Aberration-Corrected Photoacoustic Computed Tomography

### v4

@@ Abstract
Figure 1: Self-Supervised Speed of Sound Recovery Corrects PACT Aberrations. (a) Backprojection converts photoacoustic measurements into images (purple), which suffer aberrations caused by unaccounted-for variation in the speed of sound (SOS) of tissue. Our method produces high-quality images through a self-supervised recovery of SOS. We begin with a stack of aberrated reconstructions, created with a varying delay parameter (green), roughly analogous to a focal stack. Each image patch is recovered with a multi-channel deconvolution. Our forward model (dashed red) uses a trainable coordinate-based SOS representation to estimate the point spread functions (PSFs) at each location (white “x" for example shown). The physics-based forward model and multi-channel inversion are fully differentiable, enabling test-time training without any external training data, which is currently limited for this imaging modality. (b) We benchmark our method using both a neural field (NF) and pixel grid (PG) for SOS representation. Compared to the best existing methods, both with and without SOS recovery, we provide state-of-the-art image quality, and recover SOS with an order-of-magnitude speed-up.
@@ 1 Introduction
Though supervised learning methods have made significant progress on image restoration in many applications, PACT is a relatively novel imaging modality and appropriate datasets (large and naturally distributed set of measurements, with realistic aberrations, accompanied by ground truth for both image and SOS values) are not available. We consider a ring array PACT system and address these challenges with a self-supervised joint recovery method that learns SOS by backpropagation through a differentiable physical model, then uses the SOS to reconstruct a clean image. Specifically, as shown in in Fig. 1, we optimize a differentiable coordinate-based representation of the SOS, parameterized by a pixel grid (PG) or neural field (NF), to best match observed aberrations, which can then be removed with deconvolution. By imposing physical consistency across image locations and a delay-based “focal stack" (details in Sec. 3), this method converges quickly to produce an approximate SOS and high quality image. Our key contributions are:
@@ 2 Related Work
Neural fields (NFs), or implicit neural representations (INRs) have been shown to be capable of representing a variety of natural signals [26, 37], and have been demonstrated in many computational imaging tasks, including medical imaging [20, 23, 43, 28, 25, 38]. NFs represent images with a relatively small number of parameters compared to pixel grids [27], with no need for external or labeled training data. Recent works in PACT have utilized NFs for IP image reconstruction from limited or sparse measurements [46, 36, 41], but these assume constant SOS and only reconstruct the IP image. NFs have also been used to reconstruct SOS in ultrasound CT [2] and have shown promise in estimating wavefronts for computational image reconstruction [15] and closed-loop control systems for adaptive optics [7]. We, like [15], do not intervene in the measurement process, but computationally correct for wavefront-based aberration parameters estimated with a NF. We use the SIREN NF architecture, which is a fully connected network with sinusoidal activations [26].
@@ 3.1 Image Formation and Aberration
Filtered back-projection [39] is a widely used image reconstruction method in ring array PACT systems. In practice, due to the limited bandwidth and non-flat angular response of the transducers [5], a delay-and-sum (DAS) algorithm [30, 9, 11] is often used as a reasonable approximation. A delay-and-sum image $y(\mathbf{r};d)$ varies over sample locations $\mathbf{r}^{\prime}$ as well as a user-selected extra delay parameter $d$ that functions as a “focus setting" [32], and is reconstructed according to
@@
Figure 2: Acoustic heterogeneity creates non-ideal PSFs due to wavefront distortions. We consider three SOS patterns (a): the uniform SOS assumed by standard methods, a simplified body model with two internal SOS values, and a more realistic numerical phantom that we created to mimic a body slice including representative materials ranging from air to bone. For the location marked with a white “x", we illustrate (b) the discrepancy between the standard methods’ assumed wavefront and the true wavefront location arising from the SOS shown. These wavefront errors determine the PSFs (c) at the indicated location for each SOS, which lead to image aberrations for non-pointlike PSFs. A user-set delay value in the image reconstruction changes the resulting PSFs and aberrations (d), and can be used to effectively remove aberrations for simple SOS models but not realistic ones.
@@
where $\theta$ is the angle of the vector $\mathbf{r}-\mathbf{r}^{\prime}$. See Fig. 2b for example wavefront errors corresponding to the SOS maps $\mathbf{v}$ shown in Fig. 2a for the location $\mathbf{r}^{\prime}$ indicated by a white “x". It has been shown in [5] that the resulting aberrations, despite varying across the sample, can be modeled within isoplanatic image patches $y_{i}$ as convolutions with PSFs $h_{i}$ and unaberrated patches $x_{i}$:
@@ 3.2 Multi-channel Deconvolution
Figure 4: Numerical simulation demonstrates accuracy and computational efficiency. (a) Our numerical phantom allows quantitative evaluation of image reconstruction under different SOS distributions in simulation. The PA signals are generated with k-Wave simulation [31] using the IP and SOS shown. (b) The wavefront (sampled location marked by white “x" in a) reconstructed by our method with NF (purple) and PG (green) is much closer to the true wavefront (red) than other methods, better capturing image aberrations. (c) We compare our method with conventional DAS, Dual-SOS DAS, our multi-channel deconvolution with the true SOS, and APACT. While conventional DAS assumes a uniform SOS and results in significant image aberrations, Dual-SOS DAS provides better contrast by assuming a single-body SOS but cannot fully undo the aberrations and reconstruct the SOS. Given an accurate SOS, our multi-channel deconvolution is able to efficiently reconstruct a high-quality IP image. APACT is able to deblur the image effectively but does not address overall shrinkage (see vertical shift in green boxes) and cannot reconstruct an accurate SOS. Our method with NF solves the shrinking problem by fully characterizing the wavefront and is able to fully reconstruct the IP and SOS (see metrics) in a much shorter time than APACT. Ours with PG also produces good IP images but generates stripe-like artifacts in the SOS that originate from the path-integral nature of the wavefront computation. Providing our method with the ground truth SOS leads to only marginal improvements in image quality. See visualizations of the convergence in the supplementary videos.
@@ 3.3 Joint Reconstruction
We benchmark two types of SOS representation: a pixel grid (PG) with the same resolution as the IP image, and a neural field (NF) representation. For the latter, we used a single-layer, 256-feature, fully-connected network with sinusoidal activations [26]. This non-pretrained, coordinate-based network maps pixel coordinates to SOS predictions. The total number of NF parameters (1027) is much smaller than the number of pixels ($\sim$200k) in the SOS map, which reduces the degrees of freedom in the inverse problem and acts as an implicit regularizer that makes the problem better-posed. See Sec. S3.3 for ablations on network size. For both representations, we reduce unnecessary computation by restricting these pixel grid or NF coordinates to a circular mask roughly the size of the sample and fix the background SOS based on measured water temperature [19].
@@ 4.1 Numerical Simulations
The wavefront errors modeled/reconstructed (sample location marked by white “x" on SOS in Fig. 4a) by different methods are shown in Fig. 4b. To use the terminology of [5], the wavefront errors can be decomposed into a Fourier series, where the components degrade the IP image differently. Conventional DAS assumes a uniform SOS thus a constant wavefront (red), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (yellow). APACT (green) includes only the zeroth and second order Fourier series terms in its wavefront model (see [5] and Sec. S2 in the supplement for details), missing both the first order (the cosine component captured by dual-SOS in yellow) as well as higher order terms capturing finer detail. In comparison, our methods can accurately model lower and higher-order components in the wavefront (purple and blue), compared with the truth wavefront (black).

### v5

@@ Abstract
Figure 1: Self-Supervised Speed of Sound Recovery Corrects PACT Aberrations. (a) Backprojection converts photoacoustic measurements into images (purple), which suffer aberrations caused by unaccounted-for variation in the speed of sound (SOS) of tissue. Our method produces high-quality images through a self-supervised recovery of SOS. We begin with a stack of aberrated reconstructions, created with a varying delay parameter (green), roughly analogous to a focal stack. Each image patch is recovered with a multi-channel deconvolution. Our forward model (dashed red) uses a trainable coordinate-based SOS representation to estimate the point spread functions (PSFs) at each location (white “x” for example shown). The physics-based forward model and multi-channel inversion are fully differentiable, enabling test-time training without any external training data, which is currently limited for this imaging modality. (b) We benchmark our method using both a neural field (NF) and pixel grid (PG) for SOS representation. Compared to the best existing methods, both with and without SOS recovery, we provide state-of-the-art image quality, and recover SOS with an order-of-magnitude speed-up.
@@ 1 Introduction
Though supervised learning methods have made significant progress on image restoration in many applications, PACT is a relatively novel imaging modality and appropriate datasets (large and naturally distributed set of measurements, with realistic aberrations, accompanied by ground truth for both image and SOS values) are not available. We consider a ring array PACT system and address these challenges with a self-supervised joint recovery method that learns SOS by backpropagation through a differentiable physical model, then uses the SOS to reconstruct a clean image. Specifically, as shown in in Fig. 1, we optimize a differentiable coordinate-based representation of the SOS, parameterized by a pixel grid (PG) or neural field (NF), to best match observed aberrations, which can then be removed with deconvolution. By imposing physical consistency across image locations and a delay-based “focal stack” (details in Sec. 3), this method converges quickly to produce an approximate SOS and high quality image. Our key contributions are:
@@ 2 Related Work
Neural fields (NFs), or implicit neural representations (INRs) have been shown to be capable of representing a variety of natural signals [28, 38], and have been demonstrated in many computational imaging tasks, including medical imaging [22, 25, 44, 29, 27, 39]. NFs, as well as interpolated pixel grids [8], represent images with no need for external or labeled training data (see comparisons in [8, 17]). Recent works in PACT have utilized NFs for IP image reconstruction from limited or sparse measurements [47, 37, 42], but these assume constant SOS and only reconstruct the IP image. NFs have also been used to reconstruct SOS in ultrasound CT [2] and have shown promise in estimating wavefronts for computational image reconstruction [16] and closed-loop control systems for adaptive optics [7]. We, like [16], do not intervene in the measurement process, but computationally correct for wavefront-based aberration parameters estimated with a NF. We use the SIREN NF architecture, which is a fully connected network with sinusoidal activations [28].
@@ 3.1 Image Formation and Aberration
Filtered back-projection [40] is a widely used image reconstruction method in ring array PACT systems. In practice, due to the limited bandwidth and non-flat angular response of the transducers [5], a delay-and-sum (DAS) algorithm [31, 10, 12] is often used as a reasonable approximation. A delay-and-sum image $y(\mathbf{r};d)$ varies over sample locations $\mathbf{r}^{\prime}$ as well as a user-selected extra delay parameter $d$ that functions as a “focus setting” [33], and is reconstructed according to
@@
Figure 2: Acoustic heterogeneity creates non-ideal PSFs due to wavefront distortions. We consider three SOS patterns (a): the uniform SOS assumed by standard methods, a simplified body model with two internal SOS values, and a more realistic numerical phantom that we created to mimic a body slice including representative materials ranging from air to bone. For the location marked with a white “x”, we illustrate (b) the discrepancy between the standard methods’ assumed wavefront and the true wavefront location arising from the SOS shown. These wavefront errors determine the PSFs (c) at the indicated location for each SOS, which lead to image aberrations for non-pointlike PSFs. A user-set delay value in the image reconstruction changes the resulting PSFs and aberrations (d), and can be used to effectively remove aberrations for simple SOS models but not realistic ones.
@@
where $\theta$ is the angle of the vector $\mathbf{r}-\mathbf{r}^{\prime}$. See Fig. 2b for example wavefront errors corresponding to the SOS maps $\mathbf{v}$ shown in Fig. 2a for the location $\mathbf{r}^{\prime}$ indicated by a white “x”. It has been shown in [5] that the resulting aberrations, despite varying across the sample, can be modeled within isoplanatic image patches $y_{i}$ as convolutions with PSFs $h_{i}$ and unaberrated patches $x_{i}$:
@@ 3.2 Multi-channel Deconvolution
Figure 4: Numerical simulation demonstrates accuracy and computational efficiency. (a) Our numerical phantom allows quantitative evaluation of image reconstruction under different SOS distributions in simulation. The PA signals are generated with k-Wave simulation [32] using the IP and SOS shown. (b) The wavefront (sampled location marked by white “x” in a) reconstructed by our method with NF (purple) and PG (green) is much closer to the true wavefront (red) than other methods, better capturing image aberrations. (c) We compare our method with conventional DAS, Dual-SOS DAS, our multi-channel deconvolution with the true SOS, and APACT. While conventional DAS assumes a uniform SOS and results in significant image aberrations, Dual-SOS DAS provides better contrast by assuming a single-body SOS but cannot fully undo the aberrations and reconstruct the SOS. Given an accurate SOS, our multi-channel deconvolution is able to efficiently reconstruct a high-quality IP image. APACT is able to deblur the image effectively but does not address overall shrinkage (see vertical shift in green boxes) and cannot reconstruct an accurate SOS. Our method with NF solves the shrinking problem by fully characterizing the wavefront and is able to fully reconstruct the IP and SOS (see metrics) in a much shorter time than APACT. Ours with PG also produces good IP images but generates stripe-like artifacts in the SOS that originate from the path-integral nature of the wavefront computation. Providing our method with the ground truth SOS leads to only marginal improvements in image quality. See visualizations of the convergence in the supplementary videos.
@@ 3.3 Joint Reconstruction
We benchmark two types of SOS representation: an interpolated pixel grid (PG) with the same resolution as the IP image, and a neural field (NF). For the latter, we used a single-layer, 256-feature, fully-connected network with sinusoidal activations [28]. This non-pretrained, coordinate-based network maps pixel coordinates to SOS predictions. The total number of NF parameters (1027) is much smaller than the number of pixels ($\sim$200k) in the SOS map, which reduces the degrees of freedom in the inverse problem and acts as an implicit regularizer that makes the problem better-posed. See Sec. S3.3 for ablations on network size. For both representations, we reduce unnecessary computation by restricting these pixel grid or NF coordinates to a circular mask roughly the size of the sample and fix the background SOS based on measured water temperature [21].
@@ 4.1 Numerical Simulations
The wavefront errors modeled/reconstructed (sample location marked by white “x” on SOS in Fig. 4a) by different methods are shown in Fig. 4b. To use the terminology of [5], the wavefront errors can be decomposed into a Fourier series, where the components degrade the IP image differently. Conventional DAS assumes a uniform SOS thus a constant wavefront (red), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (yellow). APACT (green) includes only the zeroth and second order Fourier series terms in its wavefront model (see [5] and Sec. S2 in the supplement for details), missing both the first order (the cosine component captured by dual-SOS in yellow) as well as higher order terms capturing finer detail. In comparison, our methods can accurately model lower and higher-order components in the wavefront (purple and blue), compared with the truth wavefront (black).